OpenStudy (anonymous):

if you have cos(x) - cos(2x), is it possible to take out a cos(x) from the equation? and if so then what would that leave you left with?

myininaya (myininaya):

\[\cos(2x)=\cos^2(x)-\sin^2(x)=\cos^2(x)-(1-\cos^2(x))=2\cos^2(x)-1\] *=======================================* \[\cos(x)-\cos(2x)=\] \[\cos(x)-(2 \cos^2(x)-1)=-2\cos^2(x)+\cos(x)+1\] Are you asking if it is factorable

OpenStudy (anonymous):

well cause really the problem that i have right now is sin(2x) = cos(2x) and i need to solve for x

myininaya (myininaya):

\[\text{ Let } u=2x\] Can you solve the following for u: \[\sin(u)=\cos(u)\]

myininaya (myininaya):